solve x: sin (x+30) = - 2 cos x
1
Expert's answer
2012-11-07T09:42:24-0500
Notice that
sin(x+30) =sin(x)*cos(30) - cos(x)*sin(30)
=sqrt(3)/2 * sin(x) - 1/2 * cos(x).

Thus
sqrt(3)/2 * sin(x) - 1/2 * cos(x) = - 2 cos(x)
sqrt(3)/2 * sin(x) = -3/2 * cos(x)
sin(x)/cos(x) = -sqrt(3)
tan(x) = -sqrt(3)
x = - 60 + 180 * k, k belongs to Z